OpenStudy (konradzuse):

For series does certain summations require n = 0 or n = 1?

OpenStudy (amistre64):

it can start at any n value as long as you adjust for it

OpenStudy (konradzuse):

Or does it not matter? Originally I thought that a Geo-series must be n=0, but it seems like I've seen n=1... Would the eq then just be c^(n-1) instead of c^n?

OpenStudy (konradzuse):

Okies I guess what I was doing then....

OpenStudy (konradzuse):

Which series matters?

OpenStudy (amistre64):

spose you have the series defined by: \[\sum \frac{n(n+1)}{2}\]to define the triangle numbers 1,3,6,10, ... at n=1 this would start at 1, at n=0 it would start at 0 \[\sum_{n=1}^{inf} \frac{n(n+1)}{2}:~n=1,2,3,...\] to adjust the n, say n=0, then we adjust the formula; -1 = +1 \[\sum_{n=0}^{inf} \frac{(n+1)((n+1)+1)}{2}:~n=0,1,2,3,...\]

OpenStudy (amistre64):

if you have a series such that n=0 makes a forbidden move; like zero in a denominator, or in ln or a negative in an even root ... then adjust as needed